OpenStudy (anonymous):

x+4 -------- x2−5x+6 ÷ x2−16 ------- x+3

OpenStudy (luigi0210):

When dividing fractions, you have to switch the sign and flip the second equation: \[\LARGE \frac{x+4}{x^2-5x+6}*\frac{x+3}{x^2-16}\] Now just factor.

OpenStudy (nurali):

\[\frac{ x+4 }{ x^2-5x+6 }*\frac{ x+3 }{ x^2-16 }\] \[\frac{ x+4 }{ (x-2)(x-3) }*\frac{ x+3 }{ (x+4)(x-4)}\]

OpenStudy (nurali):

\[\frac{ x+3 }{ (x-2)(x-3)(x-4) }\]

OpenStudy (anonymous):

thank you again! @Nurali :)

OpenStudy (nurali):

My Pleasure.